# ANN-Based Grid Impedance Estimation for Adaptive Gain Scheduling in VSG Under Dynamic Grid Conditions

### v1

@@ I Introduction
The increasing penetration of renewable energy sources (RES) has led to significant changes in power-system dynamics, necessitating advanced control strategies for inverter-based resources (IBRs) [1, 2]. Unlike conventional synchronous generators, IBRs rely on power electronics to interface with the grid, which alters overall system stability characteristics. Two common control paradigms for IBRs are grid-following (GFL) inverters and grid-forming (GFM) inverters. While GFL inverters operate by tracking the grid voltage and injecting the requisite active and reactive power, they struggle under weak-grid conditions due to their dependence on external synchronization signals. In contrast, GFM inverters, particularly those based on the Virtual Synchronous Generator (VSG) concept which emulate the dynamic behavior of traditional synchronous machines, thereby enhancing system stability and providing inertia support. However, despite their advantages in weak-grid scenarios, VSGs can become unstable under strong-grid conditions if their controllers are not properly tuned.
@@ II Small signal stability issues of VSG under strong grid condition
where $A$, $B$, $C$, and $D$ are the partial derivatives of $P_{pcc}$ and $Q_{pcc}$ evaluated at the operating point. In this paper, the coupling terms $B$ and $C$ between active and reactive power are ignored when designing the controllers. Therefore, the relationships $P_{pcc}-\delta$ and $Q_{pcc}-\Delta V_{pcc}$ are presented as follow:
@@
where 00 denotes the steady-state operating point. The transfer function of active and reactive power control are shown as follow:
@@
where $D_{p}$ and $D_{q}$ are the droop coefficients, and $K_{ip}$ and $K_{iq}$ determine the virtual inertia of the VSG. From (4)–(6) and (5)–(7), the open-loop transfer functions are presented in (8) and (9), respectively.
$\begin{split}G^{P}_{op}(s)&=\frac{\Delta P_{pcc}}{\Delta P_{ref}}=\frac{K_{ip}}{s(s+D_{p}K_{ip})}\;\cdot\;\frac{3}{R_{g}^{2}+X_{g}^{2}}\\ &\quad\times\Bigl{(}R_{g}\,V_{pcc_{0}}\,V_{g}\,\sin\delta_{0}+\,X_{g}\,V_{pcc_{0}}\,V_{g}\,\cos\delta_{0}\Bigr{)}\,.\end{split}$ (8)
$\begin{split}G^{Q}_{op}(s)&=\frac{\Delta Q_{pcc}}{\Delta Q_{ref}}=\frac{K_{iq}}{s+D_{q}K_{iq}}\;\cdot\;\frac{3}{R_{g}^{2}+X_{g}^{2}}\\ &\quad\times\Bigl{(}X_{g}V_{pcc_{0}}V_{g}\sin\delta_{0}-R_{g}V_{pcc_{0}}V_{g}\cos\delta_{0}\Bigr{)}\,.\end{split}$ (9)
As shown in (8) and (9), the dynamics of the VSG depend on $D_{p}$, $D_{q}$, $K_{ip}$, $K_{iq}$, $L_{g}$, and $R_{g}$. To examine the dynamics of the VSG, the Bode diagram is plotted under different scenarios, as shown in Fig. 2. The upper plot illustrates that, with fixed VSG parameters, a stronger grid (higher SCR) reduces the phase margin, potentially causing oscillations or instability [2]. However, by adjusting the VSG parameters based on grid impedance, the phase margin can be maintained across various scenarios.
@@
However, a critical question arises: how should the VSG parameters be adjusted, and under what conditions should they be modified or maintained? To address this, an adaptive gain-scheduling control scheme is proposed, which dynamically adjusts the VSG parameters based on real-time grid impedance estimation.
@@ III-A ANN-based Online Grid Impedance Estimation
The regression plots in Fig. 5 illustrate the linear relationship between the network outputs and the true target values for the training, validation, and test sets. In each case, the points lie tightly along the ideal line ($y=x$), with correlation coefficients ($R$) approaching 1 and negligible intercepts. This alignment confirms that the model achieves near-perfect predictive accuracy across all data splits and exhibits minimal bias. Training was performed using the Levenberg–Marquardt algorithm, which combines the rapid convergence properties of the Gauss–Newton method with the stability of steepest-descent. For moderately sized feedforward networks, Levenberg–Marquardt typically requires a small number of epochs to reach a low training error, ensuring both efficiency and robust generalization. Specifically, the top plot in Fig. 6 shows the gradient during the Levenberg–Marquardt optimization process. A decreasing gradient indicates that the algorithm is approaching a critical point of the loss function. The gradient approaches zero after only nine epochs, indicating that the network has converged or is near convergence in a very short time. The middle plot displays the values of $\mu$ during training, which balances the steepest-descent and Gauss–Newton update steps. $\mu$ remains within a reasonable range, never spiking excessively, indicating a well-behaved, stable training process. The bottom plot in Fig. 6 shows the validation-check values, which increment each time the validation loss fails to improve. Here, the validation check remains at zero for all nine epochs, meaning the validation error never increased, indicating that the network either improved or maintained its generalization performance throughout. After obtaining the ANN model, it is utilized to predict new values of grid impedance in the Simulink environment.
@@
Fig. 3 illustrates the implementation of the proposed method in the Simulink environment. Due to a mismatch between the simulation sampling time ($10\ \mu\mathrm{s}$) and the ANN input sampling rate ($200\ \mu\mathrm{s}$), a rate transition block is used to resolve this issue. Since 200 points per cycle (0.02 s) must be collected before being fed to the ANN model, a buffer block accumulates input data for each estimation, introducing a one-cycle delay. After buffering, the data are normalized using the z-score method with the function:
@@ III-B Adaptive Gain-Scheduled Control Function
Based on the estimated grid impedance, the Jacobian matrix components in (3) are first recalculated. The VSG parameters are then updated as shown in Fig. 3 and expressed as follows:
@@
$D_{p}=\frac{D}{100};K_{iq}=\frac{4}{D+D_{q}},$ (11)
For the detailed calculations, see the Appendix. The proposed adaptive gain-scheduling scheme aims to maintain control performance under different grid conditions, as discussed in Section II, thereby improving VSG stability. The desired control performance targets are a settling time of $T_{s}=1\ \mathrm{s}$ and zero overshoot. However, due to the omission of active–reactive coupling terms, there may be some mismatch between the design and validation phases. This discrepancy opens research questions for future work. Further analysis is provided in the next section.
@@ IV-A Simulation scenarios
- Next 20 s: $SCR=8$ (strong grid);
@@ IV-B Simulation results and discussions
The ANN-based estimator accurately tracks actual values with minimal delay (0.02 s), significantly outperforming [7] (Fig. 9). The zoom-in plots show the tracking performance during both the steady state and the transient period of the ANN-based grid-impedance estimator. Under strong and stiff grid conditions, there are some peak errors in both $R_{g}$ and $L_{g}$ during the transient period. These errors are understandable, as SCR values of [8, 20] represent grid conditions that the ANN had not encountered before. Moreover, because the inputs do not take into account the VSG parameters, the ANN does not fully capture grid impedance dynamics during the transient period.
@@ V Conclusion
This paper presents a novel ANN-based adaptive gain-scheduling control scheme for VSGs operating under varying grid conditions. Unlike conventional fixed-gain controllers, the proposed method dynamically adjusts VSG parameters based on real-time grid-impedance estimation. Key contributions include: first, a small-signal stability analysis of the VSG under different grid conditions; second, an ANN-based online grid-impedance estimation approach with high accuracy; and third, a gain-scheduled control scheme that guarantees VSG performance across weak and very strong grid conditions. Simulation results confirm that the proposed method effectively reduces settling time and oscillations while accurately estimating grid impedance.
@@ Acknowledgment
The authors work was supported by the University of Michigan-Dearborn’s Office of Research “Research Initiation and Development”.

### v2

@@ I Introduction
The increasing penetration of renewable energy sources (RES) has led to significant changes in power-system dynamics, necessitating advanced control strategies for inverter-based resources (IBRs) [1, 2]. Unlike conventional synchronous generators, IBRs rely on power electronics to interface with the grid, which alters overall system stability characteristics. Two common control paradigms for IBRs are grid-following (GFL) inverters and grid-forming (GFM) inverters. While GFL inverters operate by tracking the grid voltage and injecting the requisite active and reactive power, they struggle under weak-grid conditions due to their dependence on external synchronization signals. In contrast, GFM inverters, particularly those based on the Virtual Synchronous Generator (VSG) concept which emulate the dynamic behavior of traditional synchronous machines, thereby enhancing system stability and providing inertia support. However, despite their advantages in weak-grid scenarios, VSGs can become unstable under strong-grid conditions if their controllers are not properly designed.
@@ II Small signal stability issues of VSG under strong grid condition
where $A$, $B$, $C$, and $D$ are the partial derivatives of $P_{pcc}$ and $Q_{pcc}$ evaluated at the operating point. In this paper, the coupling terms $B$ and $C$ between active and reactive power are ignored when designing the controllers. Therefore, the relationships $P_{pcc}-\delta$ and $Q_{pcc}-\Delta V_{pcc}$ are presented as follows:
@@
where the subscript 00 denotes the steady-state operating point. The transfer function of active and reactive power control are shown as follow:
@@
where $D_{p}$ and $D_{q}$ are the droop coefficients, and $K_{ip}$ and $K_{iq}$ determine the virtual inertia of the VSG. From (4)–(6) and (5)–(7), the open-loop transfer functions for the active and reactive power control loops are denoted as $G^{P}_{op}(s)$ and $G^{Q}_{op}(s)$, which are presented in (8) and (9), respectively.
$\begin{split}G^{P}_{op}(s)&=\frac{\Delta P_{pcc}}{\Delta P_{ref}}=\frac{K_{ip}}{s(s+D_{p}K_{ip})}\;\cdot\;\frac{3}{R_{g}^{2}+X_{g}^{2}}\\ &\cdot\Bigl{(}R_{g}\,V_{pcc_{0}}\,V_{g}\,\sin\delta_{0}+\,X_{g}\,V_{pcc_{0}}\,V_{g}\,\cos\delta_{0}\Bigr{)}\,.\end{split}$ (8)
$\begin{split}G^{Q}_{op}(s)&=\frac{\Delta Q_{pcc}}{\Delta Q_{ref}}=\frac{K_{iq}}{s+D_{q}K_{iq}}\;\cdot\;\frac{3}{R_{g}^{2}+X_{g}^{2}}\\ &\cdot\Bigl{(}X_{g}V_{pcc_{0}}V_{g}\sin\delta_{0}-R_{g}V_{pcc_{0}}V_{g}\cos\delta_{0}\Bigr{)}\,.\end{split}$ (9)
As shown in (8) and (9), the dynamics of the VSG depend on $D_{p}$, $D_{q}$, $K_{ip}$, $K_{iq}$, $L_{g}$, and $R_{g}$. To examine the dynamics of the VSG, the Bode diagram is plotted under different scenarios, as shown in Fig. 2. The upper picture illustrates that, with fixed VSG parameters, a stronger grid (higher SCR) reduces the phase margin, potentially causing oscillations or instability [2]. However, by adjusting the VSG parameters based on grid impedance, the phase margin can be maintained across various scenarios.
@@
Determining how and when to adjust VSG parameters remains a critical challenge, particularly under varying grid conditions. To address this, an adaptive gain-scheduling control scheme is proposed, which dynamically adjusts the VSG parameters based on real-time grid impedance estimation.
@@ III-A ANN-based Online Grid Impedance Estimation
The regression plots in Fig. 5 illustrate the linear relationship between the network outputs and the true target values for the training, validation, and test sets. In each case, the points lie tightly along the ideal line ($y=x$), with correlation coefficients ($R$) approaching 1 and negligible intercepts. This alignment confirms that the model achieves near-perfect predictive accuracy across all data splits and exhibits minimal bias. Training was performed using the Levenberg–Marquardt algorithm, which combines the rapid convergence properties of the Gauss–Newton method with the stability of steepest-descent [14]. For moderately sized feedforward networks, Levenberg–Marquardt typically requires a small number of epochs to reach a low training error, ensuring both efficiency and robust generalization. Specifically, the top picture in Fig. 6 shows the gradient during the Levenberg–Marquardt optimization process. A decreasing gradient indicates that the algorithm is approaching a critical point of the loss function. The gradient approaches zero after only nine epochs, indicating that the network has converged or is near convergence in a very short time. The middle plot displays the values of $\mu$ during training, which balances the steepest-descent and Gauss–Newton update steps. $\mu$ remains within a reasonable range, never spiking excessively, indicating a well-behaved, stable training process. The bottom plot in Fig. 6 shows the validation-check values, which increment each time the validation loss fails to improve. Here, the validation check remains at zero for all nine epochs, meaning the validation error never increased, indicating that the network either improved or maintained its generalization performance throughout. After obtaining the ANN model, it is utilized to predict new values of grid impedance in the Simulink environment.
@@
Fig. 3, as previous discussed, illustrates the implementation of the proposed method in the Simulink environment. Due to a mismatch between the simulation sampling time ($10\ \mu\mathrm{s}$) and the ANN input sampling rate ($200\ \mu\mathrm{s}$), a rate transition block is used to resolve this issue. Since 200 points per cycle (0.02 s) must be collected before being fed to the ANN model, a buffer block accumulates input data for each estimation, introducing a one-cycle delay. After buffering, the data are normalized using the z-score method with the function:
@@ III-B Adaptive Gain-Scheduled Control Function
Based on the estimated grid impedance, the Jacobian matrix components in (3) are first recalculated. The VSG parameters are then updated as previously shown and expressed as follows:
@@
$D_{p}=\frac{D}{100};K_{iq}=\frac{4}{D+D_{q}}.$ (11)
Detailed calculations are presented in the Appendix section. The proposed adaptive gain-scheduling scheme aims to maintain control performance under different grid conditions, as discussed in Section II, thereby improving VSG stability. The desired control performance targets are a settling time of $T_{s}=1\ \mathrm{s}$ and zero overshoot. However, due to the omission of active–reactive coupling terms, there may be some mismatch between the design and validation phases. This discrepancy opens research questions for future work. Further analysis is provided in the next section.
@@ IV-A Simulation scenarios
- Middle 20 s: $SCR=8$ (strong grid);
@@ IV-B Simulation results and discussions
The ANN-based estimator accurately tracks actual values with minimal delay (0.02 s), significantly outperforming [7] (Fig. 9). The zoom-in pictures show the tracking performance during both the steady state and the transient period of the ANN-based grid-impedance estimator. Under strong and stiff grid conditions, there are some peak errors in both $R_{g}$ and $L_{g}$ during the transient period. These errors appear because of the SCR values of [8, 20] represent grid conditions that the ANN had not encountered before. Moreover, because the inputs do not take into account the VSG parameters, the ANN does not fully capture grid impedance dynamics during the transient period.
@@ V Conclusion
This paper presented a novel ANN-based adaptive gain-scheduling control scheme for VSGs operating under varying grid conditions. Unlike conventional fixed-gain controllers, the proposed method dynamically adjusts VSG parameters based on real-time grid-impedance estimation. Key contributions include: first, a small-signal stability analysis of the VSG under different grid conditions; second, an ANN-based online grid-impedance estimation approach with high accuracy; and third, a gain-scheduled control scheme that guarantees VSG performance across weak and very strong grid conditions. Simulation results confirm that the proposed method effectively reduces settling time and oscillations while accurately estimating grid impedance.
@@ Acknowledgment
The authors work was supported by the University of Michigan-Dearborn’s Office of Research “Research Initiation and Development”. The authors have made the Matlab code and simulation files associated with this paper publicly available at: https://github.com/ManhqhUMich12/ANN_Based_GIE_AdaptiveControlGramework_VSG.